If E(X) = 5, V(X) = 3, and Y = 5 - - X, what is the mean of Y?
Answer #1

Solution :

E(Y) = 5 - E(X/3) - E(X)

= 5 - 1/3 * E(X) - E(X)

= 5 - 1/3 * 5 - 5

= -5 / 3

Mean of Y is -5/3 = -1.67